OpenStudy (anonymous):

Given cosx=-4/5 and x is in Quadrant3, find sinx

OpenStudy (anonymous):

if x is in Q3, sinx<0

OpenStudy (anonymous):

draw a triangle, that is all

Nnesha (nnesha):

 draw a triangle \[\rm sin \rm \theta = \frac{ opposite }{ hypotenuse }~~~~ \cos \theta = \frac{ adjacent }{ hypotenuse } ~~\tan \theta = \frac{ opposite }{ adjacent }\] cos =adjacent /hyp